I am having trouble solving these word problems. Can anyone help?
A washer and a dryer cost $679 combined. The washer costs $71 less than the dryer. What is the cost of the dryer?
----------------------------------------------------------------------------------
Bob will rent a car for the weekend. He can choose one of two payment plans. The first plan costs $59.98 for two days plus $.15 cents per mile. The second plan costs $47.98 for two days plus $.17 cents per mile. How many miles does Bob need to drive for the two plans to cost the same?
Thanks for your help

Answer by checkley77(12844)

X+(X-71)=679
X+X-71=679
2X=679+71
2X=750
X=750/2
X=$375 IS THE COST OF THE DRYER.
375-71=$304 IS THE COST OF THE WASHER.
PROOF:
375+304=679
679=679
-----------------------------------------------
59.98+.15X=47.98+.17X
.15X-.17X=47.98-59.98
-.02X=-12
X=-12/-.02
X=600 MILES WILL RESULT IN THE 2 PLANS BEING THE SAME.
PROOF:
59.98+.15*600=47.98+.17*600
59.98+90=47.98+102
149.98=149.98